When the first of two numbers is added to twice the second the result is 21, but when the second number is added to twice the first the result is 18. Find the two numbers.: When the first of two numbers is added to twice the second the result is 21, but when the second number is added to twice the first the result is 18. Find the two numbers.
Answer by oscargut(666)

Let a the first number and b the second one
When the first of two numbers is added to twice the second the result is 21,
then a+2b=21
when the second number is added to twice the first the result is 18.
then 2a+b =18
a+2b=21
2a+b =18
doing first eq by -2 we have
-2a+-4b=-42
2a+b=18
adding both eq
-3b=-24
b=8 and a+2(8)=21
then a+16=21
a=5
Solution: a=5 and b=8